# MuJo: Multimodal Joint Feature Space Learning for Human Activity Recognition

### v2

@@ 5.4 Multimodal Classification on MM-Fit
We evaluate the multimodal capabilities of our system using the MM-Fit dataset. Since MM-Fit comes along with videos and poses at 30 Hz, we upsample both modalities to 50 Hz. We extract HERO CLIP features from the videos following the approach outlined in Section 3. We use the sliding window method described in Section 5.2 to generate instances for pose and video inputs. We also employ the same classifier architecture $C$ but select the encoders $e_{m}$ and projections $p_{m}$ depending on the input modality. Specifically, we train individual classifier models for the pose and video input modality, using their respective encoders $e_{m}$ and projections $p_{m}$. Additionally, we experiment with multimodal input by providing all three modalities (sensor, pose, and video) to the model and concatenating the outputs from the respective projections before feeding them into the classifier head $cl$. We keep the same training setup as outlined for unimodal training in Section 5.2, perform training runs with and without the NULL class, and contrast the same three training scenarios: baseline, pre-trained, and pre-trained with frozen encoder and projection weights. For the pre-trained scenarios, we reuse the MuJo model from Section 5.2 that was pre-trained for the MM-Fit target dataset (i.e., allowing left and right wrist sensor placements).
@@
|c|c|c|ccc|ccc|
Dataset With “NULL” Method Mean Macro $F_{1}$-Score
@@
Multitask .$598\pm.377$$691\pm.445$$725\pm.040$$834\pm.055$
Table 1: Comparison of classification results on accelerometer data varying the training amount using different methods for pre-training encoders. We also investigate the effect of freezing encoders during fine-tuning.
Figure 2: Classification performance across various training data fractions. The comparison includes the baseline model and models with a pre-trained MuJo encoder and projection (both with frozen and trainable weights) on accelerometer data for all evaluated datasets.
@@ 6.1 Unimodal Classification on Real-World Accelerometer Data
Table 5.4 shows that for our method the pre-trained models with trainable weights consistently outperform the baseline with respect to Macro $F_{1}$-Score on all dataset, regardless of whether 2% or 100% of the training data is used. These findings indicate that multimodal pre-training via contrastive learning on virtually generated accelerometer data can effectively improve HAR performance on real-world sensor data. Regarding comparison with other approaches, we can see how our method is more data-efficient, as we outperform other approaches in almost all cases when 2% of the labeled training data is available. When 100% of the training data is available every method is dominant in a different dataset, but ours consistently improves on the baseline unlike other approaches. Moreover, our method presents a more consistent performance when freezing the encoder, showing the quality of the learned representation.
In general including the NULL class generally results in lower scores, but our trainable models still achieve higher scores than the baseline, even if this class was not part of our pre-training. Only for the MHEALTH dataset, when utilizing 100% of the data, the performance difference becomes marginal, with scores of 0.839 versus 0.837 without the NULL class and 0.582 versus 0.577 with the NULL class. However, when labeled data is limited, we perform significantly better than the baseline. For example, when utilizing only 2% of the train data, even the pre-trained classifiers with frozen encoder and projection layers outperform the baseline in all cases except for the MM-Fit dataset when the NULL class is included.
The boxplots presented in Figure 2 compare the baseline to classifiers pre-trained with MuJo, both with frozen and trainable encoders and projections. For various training fractions ranging from 1% to 100%, we observe that the performance of our frozen representations highly depend on the target dataset and whether the NULL class is included during training. For instance, when the NULL class is excluded, the frozen models perform comparable to the pre-trained trainable models on MYOGYM and only slightly worse on MM-Fit. On MYOGYM with the NULL class, the frozen classifiers perform comparably to the baseline, whereas in all other scenarios, they generally outperform the baseline when less training data is utilized, but perform worse when more data is included in the training. On the other hand, the pre-trained models with trainable weights constantly achieve higher scores than the baseline for all training fractions on all datasets, which demonstrates the effectiveness of our pre-trained representations for accelerometer-based activity recognition, and robustness to varying sensor placements and target domains.
We have also compared MuJo to the other pre-trained methods with different training data fractions, all with trainable encoders, see Figure 3. Here it is clear that our method is clearly the best in the MYOGYM dataset. Moreover, it outperforms all others up to 10% training fraction in all dataset other than MM-Fit, where only Multitask is better.
@@ 6.2 Multimodal Classification on MM-Fit
Leveraging multimodal inputs achieves the highest performance, suggesting that HAR benefits from the diverse perspectives of multiple modalities. In this multimodal scenario, our method, compared to the baseline, improves the Macro $F_{1}$-Score from 0.824 to 0.882 (with NULL) and from 0.897 to 0.942 (without NULL) when trained on the full dataset. Furthermore, with only 2% of the training data available, we improve performance from 0.656 to 0.734 (with NULL) and from 0.790 to 0.885 (without NULL). These findings show that the pre-trained classifiers significantly outperform the baseline and learn more efficiently with a limited amount of training data across multiple input modalities.
@@
Figure 4: Classification performance across various training data fractions on MM-Fit. The baseline model is compared to models with a pre-trained encoder and projection (both with frozen and trainable weights) on all input modalities (sensor, pose, video, and multimodal).
6.3 Zero-Shot
Table 6.3 demonstrates zero-shot results of MuJo on all four datasets. For MM-Fit and MotionSense, multimodal training outperforms unimodal training, achieving constantly higher top-1, top-3, and top-5 accuracy and Macro $F_{1}$-Scores. The MHEALTH dataset shows minimal differences between both training methods. However, for the MYOGYM dataset, which contains 30 gym exercises where all but 2 require equipment like machines, barbells, and more, mutimodal training performs worse. The reason for this might be that our pre-training was conducted on YouTube videos of fitness trainers primarily performing HIIT workouts, at home and not in the gym, focusing on bodyweight exercises. This domain shift may explain the decrease in performance from unimodal to multimodal training on MYOGYM: since the video features in our pre-training dataset lack representations of equipment-based exercises, they provide no valuable information for such types of exercises. Thus, the multimodal model might struggle more to recognize activities involving gym equipment. Consequently, unimodal training performs slightly better on this complex dataset, suggesting that target domain differences significantly impact performance.
The overall low scores are mainly due to data divergences between our pre-training data and the evaluation datasets, particularly for pose (MM-Fit) and sensor data. Since we pre-trained with virtual sensor data and poses extracted by IMUTube, our models struggle when applied to real sensor datasets and variations in pose representation (MM-Fit), leading to decreased performance. Video features perform the best out of all modalities because they were extracted in the same way during pre-training. Further, the models were not pre-trained with the NULL class, which is represented as the mean of all target class embeddings. Introducing this unknown class during evaluation increases classification complexity and leads to an additional drop in performance.
@@ 7 Conclusion and Future Work
In this study, we introduced FiMAD, a multimodal dataset that contains over 78 hours of recordings. Our method, MuJo, is a novel pre-training approach that learns a joint feature space integrating video, pose, simulated accelerometer, and textual data. Our approach provides gains in real datasets such as MM-Fit, MyoGym, MotionSense, and MHEALTH, showing that our pre-training method captures a meaningful representation for HAR. Despite being trained solely on virtual IMU data, it enhances performance on real accelerometer data, especially when less data was used for training. Evenmore, our approach offers the flexibility to be applied not only to sensor inputs but also to video and pose inputs. Finally, although our zero-shot results are not outstanding, they demonstrate that multimodal training can enhance performance compared to unimodal training, if the target domain does not differ too much from our pre-training domain.
While our study presents a robust and innovative approach to HAR by integrating multiple data modalities into a joint feature space, there are several avenues for future research and enhancement. One potential improvement is to fine-tune the text and vision models during pre-training, rather than using pre-calculated text and video embeddings. Additionally, we plan to conduct a more detailed evaluation in the future, including fine-tuning MuJo on the target dataset using real sensor data, rather than relying solely on our pre-trained model. Furthermore, we aim to enhance the handling of the NULL class, which can be achieved by using a large language model to label the chapters of FiMAD that do not contain relevant activities as NULL. This would provide a clear representation for the NULL class in our dataset. Future work also includes incorporating other proxy tasks in our framework during pre-training as well as scaling the networks. In that setup we aim to train a single MuJo model instead of training a separate one for each target dataset when the sensor positions differ.

### v3

@@ 5.3 Multimodal Classification on MM-Fit
We evaluated the multimodal capabilities of our system using the MM-Fit dataset. Since MM-Fit comes along with videos and poses at 30 Hz, we upsample both modalities to 50 Hz. We extracted HERO CLIP features from the videos following the approach outlined in Section 3. We used the sliding window method described in Section 5.2 to generate instances for pose and video inputs. We also employed the same classifier architecture $C$ but selected the encoders $e_{m}$ and projections $p_{m}$ depending on the input modality. Specifically, we trained individual classifier models for the pose and video input modality, using their respective encoders $e_{m}$ and projections $p_{m}$. Additionally, we experimented with multimodal inputs by providing all three modalities (sensor, pose, and video) to the model and concatenating the outputs from the respective projections before feeding them into the classifier head $cl$. We kept the same training setup as outlined for unimodal training in Section 5.2, performed training runs with and without the NULL class, and contrasted the same three training scenarios: baseline, pre-trained, and pre-trained with frozen encoder and projection weights. For the pre-trained scenarios, we reused the MuJo model from Section 5.2 that was pre-trained for the MM-Fit target dataset (i.e., allowing left and right wrist sensor placements).
@@
|c|c|c|ccc|ccc| Dataset With “NULL” Method Mean Macro $F_{1}$-Score
@@
Multi-task .$598\pm.377$$691\pm.445$$725\pm.040$$834\pm.055$
Table 1: Comparison of classification results on accelerometer data varying the training amount using different methods for pre-training encoders. We also investigate the effect of freezing encoders during fine-tuning. Each experiment was conducted 20 times.
@@ 6.1 Unimodal Classification on Real-World Accelerometer Data
Table 5.3 shows that for our method the pre-trained models with trainable weights consistently outperform the baseline with respect to the Macro $F_{1}$-Score on all datasets, regardless of whether 2% or 100% of the training data are used. These findings indicate that multimodal pre-training via contrastive learning on virtually generated accelerometer data can effectively improve HAR performance on real-world sensor data. In comparison with other approaches, we can see that our method is more data-efficient, as we outperform other approaches in almost all cases when 2% of the labeled training data are available. When 100% of the training data are available, every method is dominant in a different dataset, but ours consistently improves on the baseline unlike other approaches. Moreover, our method presents a more consistent performance when freezing the encoder, showing the quality of the learned representation.
In general, including the NULL class results in lower scores, but our trainable models still achieve higher scores than the baseline, even if this class was not part of our pre-training. Only for the MHEALTH dataset, when utilizing 100% of the data, the performance difference becomes marginal, with scores of 0.839 versus 0.837 without the NULL class and 0.582 versus 0.577 with the NULL class. However, when labeled data are limited, we perform significantly better than the baseline. For example, when utilizing only 2% of the train data, even the pre-trained classifiers with frozen encoder and projection layers outperform the baseline in all cases except for the MM-Fit dataset when the NULL class is included.
In Figure 2, we compare MuJo with the baseline. Moreover, we compare it with three other commonly used pre-training methods, namely Multi-task (Saeed et al., 2019), Auto-encoder (Haresamudram et al., 2019), and simCLR (Tang et al., 2021), which we also pre-trained on our FiMAD dataset. The boxplots show the Macro $F_{1}$-Scores of the experiments conducted with different training data fractions ranging from 1% to 100%, all with trainable encoders. Our method constantly achieves higher scores than the baseline for all training fractions on all datasets, which demonstrates the effectiveness of our pre-trained representations for accelerometer-based activity recognition. Moreover, it outperforms all other pre-training methods on the MyoGym dataset. Additionally, we achieve better performance for training fractions up to 10% in all dataset except MM-Fit, where only Multi-task is better. For higher train fractions, simCLR slightly outperforms us on MHEALTH, and both simCLR and Auto-encoder surpass us on MotionSense.
@@ 6.2 Multimodal Classification on MM-Fit
Leveraging multimodal inputs achieves the highest performance, suggesting that HAR benefits from the diverse perspectives of multiple modalities. In this multimodal scenario, our method, compared to the baseline, improves the Macro $F_{1}$-Score from 0.824 to 0.882 (with NULL) and from 0.897 to 0.942 (without NULL) when trained on the full dataset. Furthermore, with only 2% of the training data available, we improve performance from 0.656 to 0.734 (with NULL) and from 0.790 to 0.855 (without NULL). These findings show that the pre-trained classifiers significantly outperform the baseline and learn more efficiently with a limited amount of training data across multiple input modalities.
6.3 Zero-Shot
Table 6.3 demonstrates zero-shot results of MuJo on all four datasets. For MM-Fit and MotionSense, multimodal training outperforms unimodal training, achieving consistently higher top-1, top-3, and top-5 accuracy and Macro $F_{1}$-Scores. The MHEALTH dataset shows minimal differences between both training methods. However, for the MyoGym dataset, which contains 30 gym exercises where all but two require equipment like machines, barbells, and more, multimodal training performs worse. The reason for this might be that our pre-training was conducted on YouTube videos of fitness trainers primarily performing HIIT workouts, at home and not in the gym, focusing on bodyweight exercises. This domain shift may explain the decrease in performance from unimodal to multimodal training on MyoGym: since the video features in our pre-training dataset lack representations of equipment-based exercises, they provide no valuable information for such types of exercises. Thus, the multimodal model might struggle to recognize activities that involve gym equipment. Consequently, unimodal training performs slightly better on this complex dataset, suggesting that target domain differences significantly impact performance.
The overall low scores are mainly due to data divergences between our pre-training dataset FiMAD and the evaluation datasets, particularly for pose (MM-Fit) and sensor data. Since we pre-trained with virtual sensor data and poses extracted by IMUTube, our models struggle when applied to real sensor datasets and variations in pose representation (MM-Fit), leading to decreased performance. Video features perform the best among all modalities because they were pre-extracted with HERO CLIP in the same way for both datasets, FiMAD and MM-Fit. Further, all our models were not pre-trained with the NULL class, which is represented as the mean of all target class embeddings. Introducing this unknown class during evaluation increases classification complexity and leads to an additional drop in performance.
@@ 6.3 Zero-Shot
Figure 3: Classification performance across various training data fractions on MM-Fit, where each boxplot represents the results of 20 runs. The baseline model is compared to models with a pre-trained encoder and projection (both with frozen and trainable weights) on all input modalities (sensor, pose, video, and multimodal).
@@
{NiceTabular}
|l|rrrrrr| Modality Encoder FLOPs Encoder Params Projection FLOPs Projection Params Total FLOPs Total Params
Text 2.3601 M 1.1804 M 8.5286 M 4.2662 M 10.8887 M 5.4467 M
Pose 38.1508 M 2.3296 M 9.1840 M 4.5939 M 47.3348 M 6.9235 M
Video 655.8720 K 328.1920 K 7.2179 M 3.6109 M 7.8738 M 3.9391 M
Sensor (MyoGym) 2.1521 M 619.5140 K 7.8733 M 3.9386 M 10.0254 M 4.5581 M
Sensor (MHEALTH) 2.1917 M 619.7120 K 7.8733 M 3.9386 M 10.0650 M 4.5583 M
Sensor (MotionSense) 2.1521 M 619.5140 K 7.8733 M 3.9386 M 10.0254 M 4.5581 M
Sensor (MM-Fit) 2.1719 M 619.6130 K 7.8733 M 3.9386 M 10.0452 M 4.5582 M
Table 4: FLOPs (forward pass) and parameters for encoder and projection by modality.
6.4 Computational Requirements
Table 6.3 shows the computational requirements of our models. For each modality, we calculated the number of parameters for the encoders, projections, and their total. The table shows that all models are compact, with parameters ranging between 3.94M and 6.92M parameters, resulting in model sizes of only a few MB. Additionally, we calculated the theoretical amount of floating-point operations (FLOPs) required for a forward pass with a batch size of 1. Except for the pose encoder, all models are highly efficient, with total FLOPs below 11M. The text and video encoders are particularly compact because they operate on feature vectors pre-extracted with the ADA model for text and the HERO CLIP model for video, rather than operating on raw text and video data. For the sensor encoders, the number of sensors (one for MHEALTH, two for MM-Fit, and three for MyoGym and MotionSense) has only a minor impact on size and computational complexity. The pose encoder requires the highest number of FLOPs (38.15M), due to its complex model structure and input dimensions of (3, 17, 100), where 3 represents the number of channels, 17 the number of joints, and 100 the window size.
In addition to the inference performance of the encoders and projections, we also analyzed the time required to pre-train our MuJo model, which is notably efficient. For instance, training MuJo with all four input modalities (text, sensor, video, and pose) on a A100 GPU requires less than 1 hour and uses less than 3GB of VRAM.
7 Limitations
This study has two main limitations. First, it focuses exclusively on the fitness domain for HAR. While this specialization enables the development of a targeted and optimized approach, it restricts the generalizability of our findings to other domains, such as healthcare and general daily activity recognition. Second, the feature extraction process for each modality relies on pre-trained models. Although this leverages state-of-the-art methods and reduces computational requirements during training, it also introduces dependencies on the quality and biases of these models. For instance, pre-trained pose estimation models may perform poorly on unusual or occluded postures, and sensor-based models might not generalize well across different devices or user populations. These limitations highlight the need for more domain-agnostic and robust solutions in future work. It is also important to highlight the limited zero-shot performance when the target classes diverge too much, be it by the presence of gym equipment or of classes where the subject is not performing any exercise (null class).
@@ 8 Conclusion and Future Work
In this study, we introduced FiMAD, a multimodal dataset that contains over 78 hours of recordings. Our method, MuJo, is a novel pre-training approach that learns a joint feature space integrating video, pose, simulated accelerometer, and textual data. Our approach provides gains in real datasets such as MM-Fit, MyoGym, MotionSense, and MHEALTH, showing that our pre-training method captures a meaningful representation for HAR. Despite being trained solely on virtual IMU data, it enhances performance on real accelerometer data, especially when less data were used for training. Even more, our approach offers the flexibility to be applied not only to sensor inputs but also to video and pose inputs. Finally, although our zero-shot results are not outstanding, they demonstrate that multimodal training can enhance performance compared to unimodal training, if the target domain does not differ too much from our pre-training domain.
While our study presents a robust and innovative approach to HAR by integrating multiple data modalities into a joint feature space, there are several avenues for future research and enhancement. One potential improvement is to fine-tune the text and vision models during pre-training, rather than using pre-calculated text and video embeddings. Additionally, we plan to conduct a more detailed evaluation in the future. This will include extending our pre-training approach to other HAR data sources with real sensor data, as well as training on the target dataset itself, rather than relying solely on models pre-trained with virtual accelerometer data. Furthermore, we aim to enhance the handling of the NULL class, which can be achieved by using a large language model to label the chapters of FiMAD that do not contain relevant activities as NULL. This would provide a clear representation for the NULL class in our dataset. Future work also includes incorporating other proxy tasks in our framework during pre-training as well as scaling the networks. In that setup, we aim to train a single MuJo model instead of training a separate one for each target dataset when the sensor positions differ.